the ratio of pharmacy techs on duty at the hospital during the day vs night is 32:19. express as a percent how many pharmacy techs staff work night shift


1
Expert's answer
2021-05-18T14:21:36-0400

Percent of pharmacy techs staff work night shift =19/(32+19)x 100

=1900/51=37.25%